Calculate the heat evolved when 266 g of white
phosphorus (P4) burns in air according to the equation:

P4(s) + 5O2(g)
1
Expert's answer
2021-09-07T23:28:38-0400

Molar mass of P4 = 123.896

Moles of P4 = 266/123.896 = 2.15 moles

∆H = -3013 kJ for 1 mole

= -3013 × 2.15

= -6477.95kJ